OpenStudy (he66666):

linear algebra: cross product in r^3? is the following pairs of lines perpendicular? x= 2+2t y=-3-3t z=4+4t and x=2+t y=4-t z=5-t The answer is no but I don't know how to solve it. I think you're supposed to use the dot product or direction cosines..

OpenStudy (anonymous):

x= 2+2t y=-3-3t z=4+4t A vector parallel to this line is u={2,-3,4}

OpenStudy (anonymous):

x=2+t y=4-t z=5-t A vector parallel to this line is v={1,-1,-1}

OpenStudy (anonymous):

u dot v = 2(1) +(-3)(-1) + 4(-1)= 2 +3 -4= 1 so u and v are not perpendicular

OpenStudy (anonymous):

Since the dot product of the two parallel vectors is not zero.

OpenStudy (he66666):

Then how about this one? x=3-t y=4+4t z=2+2t and x= 2t y=3-2t z=4+2t I did the dot product and it was -6 but the answer said that it was perpendicular..

OpenStudy (anonymous):

Check that you copied the problem right.

OpenStudy (he66666):

yes I copied it right. Do you think the book is wrong?

OpenStudy (anonymous):

The first line, look at the x should that be x=3+2t?

OpenStudy (he66666):

No, it says x = 3 - t in the book. I'm assuming it's a mistake